Please help me solve the equation 2x - 6 = 2. I dont understand how to do this problem. The book isnt very explanitory. Thank YOu
Answer by tutorcecilia(2144)

2x - 6 = 2
2x-6+6=2+6
2x=+8
2x/2=8/2
x=4
.
Checking:
x=4
2x-6=2
2(4)-6=2
8-6=2
2=2 [checks out]